Shift report 20030304 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 04 2003 - 16:09:33 EST

  • Next message: Flemming Videbaek: "Tuesday eve/night beam exp."
    ------------------------------------------------------------
    Date: Tue Mar  4 16:09:32 2003, Shift: 08:00-16:00
    Supervisor: CEJ, with: ZY
    Log book#14 pages 45 - 45
    
    Shift Summary:
    ------------------------------------------------------------
    1) Short store with high background in the beginning of the shift. No beam untill 14.30. Good store with low backgorund rates (compared to last night).
    
    2) Run 8199 & 8200 (high backgound, no good)
       Run 8201 MRS@40(A1050), FS@4(B1/2)(135 K trigger 2).
       Run 8202 (running...)
    
    3) Still problems with T3 (some HV channels were turned on during run 8201). Else no problems reported.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 04 2003 - 16:11:05 EST